OpenStudy (anonymous):

(6a + b) ^2 + (product)

OpenStudy (anonymous):

This is a distributive problem in mathematics. So you do the first term, then second term in the parenthesis normally. Since it has a square (or to the power of two) you do this: (6a+b)(6a+b) First term 6a x 6a = 36a^2 6a x b = 6ab Now second term b x 6a = 6ab b x b = b^2 simplify terms: 36a^2 + b^2 + 12ab